Technical Staff Blog

Author archives: Phirum Peang

RSS feed of Phirum Peang

Last update on .

The CIS Networking Group will be swapping out a UPS in the 5th floor networking closet on Tuesday, Oct. 25th, 6am-7am. The networking switches in the closet will loose power during this maintenance.  These switches provide networking connections to southwest section of the CIT building on the 4th and 5th.

The following rooms on the ...

Last update on .

All tstaff managed linux desktop will be rebooted on Wednesday, August 31st, at 7am.

Tstaff is pushing out a new nvidia driver for all managed linux desktop running X in order support new nvidia graphics hardware.  Any managed linux desktop with the following nvidia cards listed below will get the new 367.44 driver.  A ...

Last update on .

CIS Networking Group announced CS edge switches maintenance schedule on August 2nd, 6am - 8am.  See below.   

During this maintenance windows, all desktop machine network connection will be affected.  This includes most of ssh forwarding hosts except a few virtualize hosts like unison, oolong, kui, emma, eternity, earl and bushtea. A few servers will also be ...

Last update on .

We had a quick 2 minute GPFS departmental cluster nfs blip at 8:27am when one of the cluster nfs server, nerds, experienced a kernel panic.  The failover of the nfs ip address to its partner was success.  We are debugging the issue with nerds and will bring it back into the GPFS departmental nfs ...

Last update on .

CIS Networking group is planning to make changes to the 10G switches, cit-cs-r14-sw-1 and cit-cs-r14-sw-2.  The switch config change requires a reboot.  They will be changing and rebooting the cit-cs-r14-sw-1 on Tuesday, June 21st at 5:30am and cit-cs-r14-sw-2 on Thursday, June 23rd at 6:00am.  The GPFS NSD nodes are connected to these switches ...

Last update on .

We had an emergency failover of the CS firewall at 4:45pm because two security updates screwed up the active firewall software configurations.  The fail over was a manually process because some of active firewall software were removed due to a package conflict with the security updates.  There was about a ten minutes outage for ...

Last update on .

On Saturday, April 16th, the CIS Networking Group will be completing the upgrade of their two core routers.  The first upgrade is scheduled for 6am and the second will be 7am.  There might be brief connectivity problems during the maintenance windows.  Tstaff will be monitoring and verifying that our network connections are working properly after ...

Last update on .

CIS Networking Group added a new DNS server to their proxy this morning.  It's not clear how yet, but this change prevented our DNS servers from resolving DNS records external to the department. DNS records have a certain lifetime, after which they timeout. As more records timed out, our backlog became increasingly worse. Once ...

Last update on .

On Wednesday, March 30th starting at 7am, we will be restarting the departmental NFS and CIFS cluster to flush their DNS caches. 

For linux users, each NFS server restart can potentially cause a brief 90 second NFS blip for departmental linux machines. Please do not leave any unsaved files open during this time.

For Windows ...

Last update on .

There was about a 15 minute CIFS and NFS file server blip at 10:05am due to dns service problem.  We were decommissioning the old internal dns servers when this happened.  The new virtualized internal dns servers were up and running fine but a misconfigured dns acl prevented them from being use.  All CIFS and ...